OpenStudy (anonymous):

ln6 - ln2 equals what in terms of ln? ln6 - ln2 equals what in terms of ln? @Mathematics

OpenStudy (anonymous):

Do you remember the laws of logarithms?

OpenStudy (kainui):

ln(a)-ln(b) = ln(a/b)

myininaya (myininaya):

or use ln(xy)=ln(x)+ln(y) ln(6)=ln(2*3)=ln(2)+ln(3) now what is the following simplified: ln(2)+ln(3)-ln(2)

myininaya (myininaya):

but using the other law that kainui works too

OpenStudy (anonymous):

thanx everyone--it's been forever since i did this stuff!

OpenStudy (anonymous):

ln 3

myininaya (myininaya):

yep yep

OpenStudy (anonymous):

Does anyone have a copy of the rules of ln?